if the length of the rectangle is 4 more than three times the width and the area is equal to 39^2. find the dimension of the rectangle
Answer by dkppathak(439)

if the length of the rectangle is 4 more than three times the width and the area is equal to 39^2. find the dimension of the rectangle
solution
Let width is x mtrs than length =3x+4
area= lengthx width
39=x(3x+4)
3x^2+4x-39=0
3x^2+13x-9x-39=0
(3x+13)(x-3)=0
x=3 or x=-13/3 rejecting negative value of x therefore x=3 means width is 3 mtrs and length will be 13 mtrs
answer Length=13 mtrs
width=3 mtrs